OpenStudy (anonymous):

What is the average (arithmetic mean) of x, y, and z? If x+y=10; y+z=15; x+z=17.

OpenStudy (aravindg):

so we are looking for \[\large \dfrac{x+y+z}{2}\]

OpenStudy (aravindg):

Can you try to get to this form ?

OpenStudy (aravindg):

just give it a try using the given equations

OpenStudy (amoodarya):

compute sum of them then you have the answer

OpenStudy (aravindg):

@Elak any idea?

OpenStudy (amoodarya):

x+y+y+z+z+x=10+17+15 so its easy to go on!

OpenStudy (aravindg):

@amoodarya you just spoiled the questions catch :(

OpenStudy (anonymous):

is it 8?

OpenStudy (anonymous):

2x+2y+2z=48 x+y+z=48/6 ???

OpenStudy (amoodarya):

(x+y+z)/3=(10+15+17)/3=42/3=14

OpenStudy (anonymous):

:D shame on me :(

OpenStudy (amoodarya):

other way to compute is solve the system of equation and then calculate mean value

OpenStudy (anonymous):

thank you very much AravindG and amoodarya

OpenStudy (amoodarya):

x+y=10 y+z=15 z+x=17 sum of three eq. is 2(x+y+z)=10+15+17=42 so x+y+z=21 now compare it with first equation

OpenStudy (anonymous):

yes, it so easy )))

OpenStudy (aravindg):

yw :) and \[\huge \color{green}{\text{WELCOME TO OPENSTUDY!}}\]